OpenStudy (anonymous):

A. A series RC circuit has a resistance of 150 ohms and a capacitance of 40 microfarads. The capacitor is to be charged by a 120-V, 60-Hz emf source. At 1.2 milliseconds thorugh the charging process, determine the following. 1.) the capacitive time constant, 2.) the charge in the capacitor, 3.) the energy stored in the capacitor.

OpenStudy (anonymous):

\[\tau = R*C \] you can use this for 1 For 2, Q=C*V \[V(t)=V _{0} (1-e ^{\frac{ t }{ \tau }}) \] you should take V from first equation and put it into second equation. For 3, \[E=\frac{ C*V ^{2} }{ 2 } \]